# ConGeo: Robust Cross-view Geo-localization across Ground View Variations

### v1

@@ 5.1 Performance under Different Settings
As in previous works [2], we assess the effectiveness of ConGeo in the unknown orientation setting and the limited FoV settings on the CVUSA and CVACT validation datasets in Table 1. Unlike previous methods [2, 30, 12], which train different models for each setting to improve performance, the single model trained with ConGeo excels in all settings. As shown in the rows highlighted in grey in Table 1, ConGeo improves the base model’s (Sample4Geo) performance by a large margin (68.9% and 53.4% R@1 improvement for ${}^{\circ}$ and ${}^{\circ}$, respectively) and significantly outperforms the state-of-the-art methods on most of the challenging settings on the CVUSA dataset, without training models on separate FoVs. Besides, when training ConGeo on each FoV separately, the model’s performance can be further enhanced in different settings, significantly surpassing all competitors on all FoVs.
@@
Similarly, ConGeo outperforms the state-of-the-art model’s R@1 under ${}^{\circ}$, ${}^{\circ}$, ${}^{\circ}$, and ${}^{\circ}$ settings on the CVACT validation set by 0.6%, 11.4%, 12.7%, and 5.8% respectively. ConGeo also shines on this dataset with a single model, surpassing the previous approaches in most settings by a substantial margin. In short, these results show the consistent superiority of ConGeo over previous state-of-the-art methods and its robustness in handling ground view variations.
@@
Methods CVUSA CVACT Val CVACT Test R@1 R@1% R@1 R@1% R@1 R@1% CVFT [29] 61.4 99.0 61.1 95.9 - - DSM [2] 92.0 99.7 82.5 97.3 - - TransGeo [4] 94.1 99.8 85.0 98.4 - - GeoDTR [31] 95.4 99.9 86.2 98.8 64.5 98.7 SAIG-D [12] 96.3 99.9 89.1 98.9 67.5 96.8 Sample4Geo [5] 98.7 99.9 90.8 98.8 71.5 98.7 ConGeo 98.3 99.9 90.1 98.2 71.7 98.3 Table 2: Comparison of the North-aligned setting on CVUSA and CVACT datasets. The second-best performance is underlined. “-” means the score is not provided in the original paper.
@@
Figure 3: Examples of the top-4 retrieved images when ${}^{\circ}$ of ConGeo and the baseline. Images in the orange box denote the correct results.
@@
Moreover, ConGeo maintains competitive performance in the North-aligned setting. As shown in Table 2, ConGeo always ranks within the top two for R@1 both on the CVUSA and CVACT datasets. Moreover, by comparing Table 2 and Table 1, we observe that when confronting challenging settings, baselines experience a considerable performance drop (as seen from the results with grey background in both tables). For example, for a query image with ${}^{\circ}$, R@1 of the base model drops by 82.4%; while shifting the North-aligned ground view with unknown orientation leads to an R@1 decrease of 94.6%. On the contrary, ConGeo maintains robustness to various ground view shifts with R@1 dropping by only 12.8% for ${}^{\circ}$ and improving by 5.7% for ${}^{\circ}$.
@@
Fig. 3 shows two examples of top-4 retrieved images when the query image’s ${}^{\circ}$. Limited FoV images offer restricted semantic and spatial information and thus are more challenging. In Fig. 3 (a), the baseline model only retrieves aerial images with a horizontal road, indicating that the model heavily relies on spatial correspondence; ConGeo’s results are more diverse among road directions. Fig. 3 (b) shows a building next to a half-circle side road. Several images retrieved by ConGeo include similar visual features, showing that ConGeo elicits the model’s focus on information that is consistent across ground view variations.
@@ 5.2 Ablations
Comparison with Data Augmentations (Table 3). One possible way to improve robustness under view variations is through data augmentation. We compare ConGeo with three methods: “Shift” (random cyclical shift), “FoV” (random query image FoV cropping, from ${}^{\circ}$ to ${}^{\circ}$), and “Rotate” (random aerial images rotation by ${}^{\circ}$, ${}^{\circ}$ or ${}^{\circ}$). Results show that these methods can improve the base model under unknown orientation and limited FoV. However, their overall performance stays far lower than ConGeo, especially since their performance in the North-aligned setting notably decreases, showing the superiority of contrastive objectives over direct data augmentation. Compared to ConGeo, all the data augmentation approaches lack a crucial single-modal contrastive objective, which explicitly enforces discovering the joint features shared by different ground view variations and therefore being robust to them.
Aug. type North-aligned ${}^{\circ}$ ${}^{\circ}$ ${}^{\circ}$ Avg.
@@
Table 3: The comparison of results between ConGeo and task-specific data augmentations. We incorporate augmentations into Sample4Geo, “Shift” denotes using shifted query images and “FoV” denotes using query images of limited FoVs, “Rotate” randomly rotating aerial images with an angle in {${}^{\circ}$, ${}^{\circ}$, ${}^{\circ}$} as data augmentation. The second-best performance is underlined.
Ablations of Loss Components (Table 5). The aerial view single-modal contrastive objective ($\mathcal{L}_{\text{single-r}}$) slightly enhances the performance (Row 1 and 2), which serves as the basis for the following ablations. The ground view single-modal contrastive loss ($\mathcal{L}_{\text{single-q}}$) gradually boosts performance under FoVs as contrastive targets between ground views are added (Row 1, 3, and 4). Finally, adding the cross-modal contrastive loss on top of single-modal losses yields a notable improvement (Row 4 and 7). This indicates that the contrastive learning between the original and shifted ground images plays an essential role in assisting cross-modal alignment.
Table 4: Ablation studies on ${}^{\circ}$ and ${}^{\circ}$ on the CVUSA dataset. “Shift” and “FoV” mean cyclic shift and FoV cropping for ground view images.
No. $\mathcal{L}_{\text{single-r}}$ $\mathcal{L}_{\text{single-q}}$ $\mathcal{L}_{\text{cross}}$ ${}^{\circ}$ ${}^{\circ}$
@@
1 2.5 26.7 1.5 20.4
2 7.7 43.3 5.8 38.9
3 17.8 73.0 12.0 56.3
4 20.5 59.9 14.2 53.4
5 40.2 89.0 29.6 76.0
6 35.8 81.6 20.3 67.6
7 55.9 90.9 37.0 81.4
Methods ${}^{\circ}$ ${}^{\circ}$ ${}^{\circ}$
@@
Table 4: Ablation studies on ${}^{\circ}$ and ${}^{\circ}$ on the CVUSA dataset. “Shift” and “FoV” mean cyclic shift and FoV cropping for ground view images.
@@ 5.3 Adaptability to Different Base Models
ConGeo is model-agnostic: it can be plugged into different cross-view geo-localization systems and boost their robustness to ground view variations. To demonstrate this, we choose two representative methods, TransGeo [4] and SAIG-D [12] as base models, on top of Sample4Geo. For each base model, we follow the default configurations — including backbone and data augmentations — and use the vanilla loss in the base model instead of Eq.(5). We train the ConGeo-augmented models on all FoVs jointly (same as the results with grey background in Table 1). We compare each model with two baselines: (1) the base model trained on North-aligned data, (2) the base model trained with strong data augmentations, including shifting each ground image with a random angle and cropping it with a FoV between ${}^{\circ}$ and ${}^{\circ}$ (referred to as “DA”). As shown in Table 5, adding ConGeo improves ViT-based TranGeo’s R@1% performance by 37.8% and 71.1% for ${}^{\circ}$ and ${}^{\circ}$, respectively. SAIG-D combines convolutional stem and self-attention layers in the encoder and the improvement with ConGeo is also remarkable. ConGeo-augmented models also outperform strong data augmentations under most FoVs. In summary, ConGeo can be plugged into three different baselines with both CNN-based and ViT-based backbones, demonstrating its versatility and effectiveness across model architectures.
@@ 5.4 Robustness across Other Ground View Variations
Robustness across Locations. In cross-view geo-localization, the VIGOR dataset is particularly challenging, because the two views are not center-aligned. In particular, VIGOR’s cross-area subset is regarded as a standard benchmark to test the model’s robustness to data across different locations, since training and test images are taken from different areas [11, 32, 5]. As shown in Table 7, ConGeo outperforms the baselines on both the cross-area and same-area subsets. Note that we report the performance of a single model for all FoVs, for ConGeo and Sample4Geo (grey background). The results indicate that ConGeo consistently improves the model’s robustness on cross-area data.
@@
Table 7: Comparison on University-1652 [28]. ${}^{\star}$ denotes the reproduced model since the pre-trained weights are not provided.
Methods Random FoVs Random Zooming Gaussian Noise Motion Blur
R@1 R@5 R@10 R@1% R@1 R@5 R@10 R@1% R@1 R@5 R@10 R@1% R@1 R@5 R@10 R@1%
Sample4Geo 15.7 25.3 31.2 51.7 48.2 60.8 65.8 80.1 42.4 60.9 67.0 81.6 31.4 38.6 42.0 52.9
Sample4Geo + DA 85.5 95.1 96.6 98.8 44.5 60.6 66.7 82.7 0.2 1.0 1.7 8.8 16.8 23.2 25.9 36.0
ConGeo 84.2 95.1 97.0 99.4 68.7 80.3 83.8 92.1 45.8 64.4 70.3 83.4 32.5 40.7 43.5 53.5
Table 8: Comparison on unseen ground view variations between ConGeo and baselines on the CVUSA dataset. Random FoVs are between ${}^{\circ}$ to ${}^{\circ}$, Random Zooming is performed with a ratio between 0.5 and 2.0, Gaussian Noise [34], and Motion Blur [34] are added with severity: 5.
Robustness to Street Images. Street images in University-1652 are typical examples of real-world limited FoV images. The St2S and S2St settings pose significant challenges as the limited FoV images can be captured from diverse locations with unknown orientations. Differently from the experiments considering North-aligned panoramas, here we randomly sample two street view images from one location as the inputs to conduct single-modal contrastive loss. Detailed descriptions can be found in the supplementary materials. Results in Table 7 show that the performance on both settings is improved.
Robustness to Unseen Ground View Variations. In real-world scenarios, the variations of the ground view image are more diverse than orientation and FoV variations. In Table 8, we evaluate the model across four unseen ground view variations on the CVUSA dataset: Random FoVs, Random Zooming, Gaussian Noise, and Motion Blur. The ConGeo-augmented model exhibits comparable performance to the model with data augmentation on Random FoVs test, while demonstrating significant advantages over the two baselines when tested on other unseen ground view variations. As expected, task-oriented data augmentation shows unsatisfactory robustness over unseen variations, further verifying the effectiveness and generalizability of ConGeo by learning robust features from ground view variants.
@@
Figure 5: ConGeo’s activation areas are more consistent across ground view shits. The Grad-CAM activation maps of the base model and ConGeo on the North-aligned and the unknown orientation setting. The orange box indicates the same area in different ground view variants.
Figure 6: ConGeo focuses on discriminant features for the task and its activated regions are consistent across view variation. Left: GradCam activations of the base model and ConGeo (green background) on the North-aligned and FoV = ${}^{\circ}$ settings. Right: focus on the regions highlighted by the coloured boxes.
@@ 7 Limitations
We showed that ConGeo leads to significant improvements when facing arbitrarily oriented ground images and diverse FoVs, except in the North-aligned setting. This performance drop is nearly unavoidable, as ConGeo tends to disrupt the over-reliance on spatial correspondence shortcuts. However, by keeping the original learning objectives of the base model (vanilla loss), ConGeo achieves competitive performance when orientation information is available (Table 2), and significantly improves the robustness when it is unknown. Additionally, we focus on ground view orientation and FoV variations, but other variations can be envisaged (e.g., zoom, color intensity, blur). We show ConGeo’s robustness to some of these variations unseen during training (Table 8). Finally, other aspects of the model can face robustness issues and could be discussed, such as network structure and sample information; e.g., some samples with less salient semantic information might be more sensitive to ground view variations.
@@ 8 Conclusion
Tackling situations where the ground view image orientation is unknown, or the FoV is limited, is crucial for real-world applications of cross-view image geo-localization. We propose ConGeo, a single- and cross-modal contrastive method that enhances a model’s robustness to ground view variations by aligning image variations with their original representations. Experiments on four datasets and activation map analysis demonstrate that ConGeo consistently boosts the performance of state-of-the-art methods by a large margin when facing diverse challenging settings. With its adaptability to different base models and versatility to accommodate diverse ground view variations, ConGeo strives to be a step towards widening the applicability of geo-localization methods to the real world.
@@ Acknowledgments
We thank Haoyuan Li, Zimin Xia, Gencer Sumbul, Silin Gao, Valérie Zermatten, Zeming Chen, Tianqing Fang, Gaston Lenczner, Kuangyi Chen, Robin Zbinden, Giacomo May, Riccardo Ricci, Emanuele Dalsasso, and Sepideh Mamooler for providing helpful feedback on earlier versions of this work. We acknowledge the support from the EPFL Science Seed Fund. AB gratefully acknowledges the support of the Swiss National Science Foundation (No. 215390), Innosuisse (PFFS-21-29), the EPFL Center for Imaging, Sony Group Corporation, and the Allen Institute for AI.
Supplementary Materials
@@
  • Additional experiments: architecture analysis (CNN-based and ViT-based encoders), transfer experiments, polar transformation (Section S5).
@@
  • Visualization: more samples and failure case analysis (Section S7).
@@
S3 Detailed Method Description
@@
Training phase: In each mini-batch, we take paired ground-aerial images ($I_{Q}$, $I_{R}$) as input, and perform defined transformations $T_{q}$ (i.e., random shift, FoV cropping) to each ground view image $I_{q}$ ($I_{q}\in I_{Q}$), obtaining the transformed ground image $I_{q}{{}^{*}}=T_{q}(I_{q}|\theta,\alpha)$. We also perform augmentations in the base model to $I_{R}$ (described in the data preprocessing of Section S2), obtaining $*I_{R}{{}^{*}}$. Then, the image batchs $I_{Q}$, $*I_{Q}{{}^{*}}$, $I_{R}$ and $*I_{R}{{}^{*}}$ are encoded by two encoders, a query encoder $E_{q}$ for encoding $I_{Q}$, $*I_{Q}{{}^{*}}$, and a reference encoder $E_{r}$ for encoding $I_{R}$, $*I_{R}{{}^{*}}$, with output feature embeddings $Q$, $*Q{{}^{*}}$, $R$, and $*R{{}^{*}}$ respectively. The optimization goal contains four targets, two single-modal contrastive objectives, one for aligning the feature space between $Q$ and $*Q{{}^{*}}$ while another for aligning the feature space between $R$ and $*R{{}^{*}}$, a cross-modal contrastive objective aligning the feature space between $R$ and $*Q{{}^{*}}$, and a vanilla contrastive objective aligning the feature space between $R$ and $Q$. The total loss is a weighted summation of the four loss functions, where the ablation of weights will be studied in the next section.
@@
Street images in the University-1652 dataset are ground view images with unknown orientations and limited FoVs. In order to adapt ConGeo to street images without collecting a full-view panorama, we provide an alternative way of building the ground view single-modal contrastive objective (Eq. (1) in the paper). Instead of enforcing the representations of ground view variants to be closer to the original ones, we construct the single-modal contrastive objective by emphasizing the proximity between street images obtained from the same geographic locations but depicting different perspectives or angles. Since the street-to-satellite retrieval is a many-to-one matching, there is more than one street view image located in the region of satellite images. Therefore, for each street view sample, we randomly select another street image with the same location to construct the single-modal contrastive objective for training.
S4 Hyper-parameter Analysis
@@
In order to achieve robustness across ground view variations, ConGeo applies a set of transformations on ground view images during training. The transformation includes shifting the ground view images with a random orientation angle $\theta$ and cropping the shifted images with a FoV angle $\alpha$. In the main experiments, we empirically set $\alpha$ to ${}^{\circ}$. Here, we provide detailed results by using different training FoV angles. Results are shown in Table S2. In the table, we report models trained with FoV images from ${}^{\circ}$, ${}^{\circ}$, ${}^{\circ}$, and random FoV from ${}^{\circ}$ to ${}^{\circ}$ on the unknown orientation and FoV settings. It can be seen that ConGeo’s performance is robust under different training FoVs, which consistently surpasses the baseline with targeted data augmentation by a large margin.
Method($\alpha$) ${}^{\circ}$ ${}^{\circ}$ ${}^{\circ}$ ${}^{\circ}$
R@1 R@5 R@10 R@1% R@1 R@5 R@10 R@1% R@1 R@5 R@10 R@1% R@1 R@5 R@10 R@1%
Sample4Geo+DA 84.1 95.0 96.9 99.4 63.6 84.6 90.1 98.2 32.2 55.1 64.5 87.4 21.5 42.0 51.6 79.9
ConGeo(180) 85.2 95.1 96.9 98.9 92.3 97.9 98.7 99.7 55.9 73.2 79.0 90.9 37.1 55.7 62.8 81.4
ConGeo(270) 70.4 86.6 90.8 97.4 77.9 91.7 94.2 98.2 44.9 66.1 73.0 88.1 28.1 48.9 57.3 79.6
ConGeo(360) 96.6 98.9 99.2 99.7 83.8 94.2 96.1 98.8 38.2 58.2 65.2 82.0 19.5 35.9 43.8 66.5
ConGeo(180-360) 83.1 93.5 95.7 98.4 82.9 93.9 95.8 98.8 31.0 50.1 58.0 77.2 28.9 49.4 57.5 78.4
Table S2: Analysis of training ConGeo with different FoV angle $\alpha$: ${}^{\circ}$, ${}^{\circ}$, ${}^{\circ}$, and random angles between ${}^{\circ}$${}^{\circ}$, on the CVUSA dataset. Sample4Geo+DA means training Sample4Geo with targeted data augmentation (random shift and random FoVs ${}^{\circ}$${}^{\circ}$).
Table S3: Analysis of loss weights of single- and cross-modal contrastive objectives on the CVUSA dataset.
Loss weights North-aligned ${}^{\circ}$
@@
LR North-aligned ${}^{\circ}$
@@
Table S3: Analysis of loss weights of single- and cross-modal contrastive objectives on the CVUSA dataset.

### v2

@@ 5.1 Performance under Different Settings
As in previous works [27], we assess the effectiveness of ConGeo in the unknown orientation setting and the limited FoV settings on the CVUSA and CVACT validation datasets in Table 1. Unlike previous methods [27, 36, 45], which train different models for each setting to improve performance, the single model trained with ConGeo excels in all settings. As shown in the rows highlighted in grey in Table 1, ConGeo improves the base model’s (Sample4Geo) performance by a large margin (68.9% and 53.4% R@1 improvement for FoV=360 and FoV=90, respectively) and significantly outperforms the state-of-the-art methods on most of the challenging settings on the CVUSA dataset, without training models on separate FoVs. Besides, when training ConGeo on each FoV separately, the model’s performance can be further enhanced in different settings, significantly surpassing all competitors on all FoVs.
@@
Similarly, ConGeo outperforms the state-of-the-art model’s R@1 under 360, 180, 90, and 70 settings on the CVACT validation set by 0.6%, 11.4%, 12.7%, and 5.8% respectively. ConGeo also shines on this dataset with a single model, surpassing the previous approaches in most settings by a substantial margin. In short, these results show the consistent superiority of ConGeo over previous state-of-the-art methods and its robustness in handling ground view variations.
Methods CVUSA CVACT Val CVACT Test
R@1 R@1% R@1 R@1% R@1 R@1%
CVFT [28] 61.4 99.0 61.1 95.9 - -
DSM [27] 92.0 99.7 82.5 97.3 - -
TransGeo [42] 94.1 99.8 85.0 98.4 - -
GeoDTR [40] 95.4 99.9 86.2 98.8 64.5 98.7
SAIG-D [45] 96.3 99.9 89.1 98.9 67.5 96.8
Sample4Geo [5] 98.7 99.9 90.8 98.8 71.5 98.7
ConGeo 98.3 99.9 90.1 98.2 71.7 98.3
@@
Table 2: Comparison of the North-aligned setting on CVUSA and CVACT datasets. The second-best performance is underlined. “-” means the score is not provided in the original paper.
@@
Figure 3: Examples of the top-4 retrieved images from ConGeo and the baseline when FoV=90. Images in the orange box denote the correct results.
@@
Moreover, ConGeo maintains competitive performance in the North-aligned setting. As shown in Table 2, ConGeo always ranks within the top two for R@1 both on the CVUSA and CVACT datasets. Moreover, by comparing Table 2 and Table 1, we observe that when confronting challenging settings, baselines experience a considerable performance drop (as seen from the results with grey background in both tables). For example, for a query image with FoV=180, R@1 of the base model drops by 82.4%; while shifting the North-aligned ground view with unknown orientation leads to an R@1 decrease of 94.6%. On the contrary, ConGeo maintains robustness to various ground view shifts with R@1 dropping by only 13.1% for 360 and by 6.0% for 180.
@@
Fig. 3 shows two examples of top-4 retrieved images when the query image’s FoV=90. Limited FoV images offer restricted semantic and spatial information and thus are more challenging. In Fig. 3 (a), the baseline model only retrieves aerial images with a horizontal road, indicating that the model heavily relies on spatial correspondence; ConGeo’s results are more diverse among road directions. Fig. 3 (b) shows a building next to a half-circle side road. Several images retrieved by ConGeo include similar visual features, showing that ConGeo elicits the model’s focus on information that is consistent across ground view variations.
@@ 5.2 Ablations
Comparison with Data Augmentations (Table 3). One possible way to improve robustness under view variations is through data augmentation. We compare ConGeo with three methods: “Shift” (random cyclical shift), “FoV” (random query image FoV cropping, from 70 to 360), and “Rotate” (random aerial images rotation by 90, 180 or 270). Results show that these methods can improve the base model under unknown orientation and limited FoV. However, their overall performance stays far lower than ConGeo, especially since their performance in the North-aligned setting notably decreases, showing the superiority of contrastive objectives over targeted data augmentation. Compared to ConGeo, all the data augmentation approaches lack a crucial single-view contrastive objective, which explicitly enforces discovering the joint features shared by different ground view variations and therefore being robust to them.
Aug. type North-aligned FoV=360 FoV=180 FoV=90 Avg.
@@
Table 3: The comparison of results between ConGeo and task-specific data augmentations. We incorporate augmentations into Sample4Geo, “Shift” denotes using shifted query images and “FoV” denotes using query images of limited FoVs, “Rotate” randomly rotating aerial images with an angle in {90, 180, 270} as data augmentation. The second-best performance is underlined.
Ablations of Loss Components (Table 5). The aerial view contrastive objective ($\mathcal{L}_{\text{single-r}}$) slightly enhances the performance (Row 1 and 2), which serves as the basis for the following ablations. The ground view contrastive loss ($\mathcal{L}_{\text{single-q}}$) gradually boosts performance under FoVs as contrastive targets between ground views are added (Row 1, 3, and 4). Finally, adding the cross-view contrastive loss on top of single-view losses yields a notable improvement (Row 4 and 7). This indicates that the contrastive learning between the original and shifted ground images plays an essential role in assisting cross-view alignment.
No. $\mathcal{L}_{\text{single-r}}$ $\mathcal{L}_{\text{single-q}}$ $\mathcal{L}_{\text{cross}}$ FoV=180 FoV=90
@@
1 4.1 30.4 2.5 26.7
2 15.7 60.0 7.7 43.3
3 44.1 93.4 17.8 73.0
4 37.9 81.2 20.5 59.9
5 91.5 99.7 40.2 89.0
6 81.7 98.8 35.8 81.6
7 92.3 99.7 55.9 90.9
Table 4: Ablation studies on FoV=180 and FoV=90 on the CVUSA dataset. “Shift” and “FoV” mean cyclic shift and FoV cropping for ground view images.
Methods FoV=360 FoV=180 FoV=90
@@ 5.3 Adaptability to Different Base Models
ConGeo is model-agnostic: it can be plugged into different CVGL systems and boosts their robustness to ground view variations. To demonstrate this, we choose two representative methods, TransGeo [42] and SAIG-D [45] as base models, besides Sample4Geo. For each base model, we follow the default configurations — including backbone and data augmentations — and use the vanilla loss in the base model instead of Eq.(5). We train the ConGeo-augmented models on all FoVs jointly (same as the results with grey background in Table 1). We compare each model with two baselines: (1) the base model trained on North-aligned data, (2) the base model trained with strong data augmentations, including shifting each ground image with a random angle and cropping it with a FoV between 70 and 360 (referred to as “DA”). As shown in Table 5, adding ConGeo improves ViT-based TranGeo’s R@1% performance by 37.8% and 69.9% for FoV=360 and FoV=90, respectively. SAIG-D combines convolutional stem and self-attention layers in the encoder and the improvement with ConGeo is also remarkable. ConGeo-augmented models also outperform strong data augmentations under most FoVs. In summary, ConGeo can be plugged into three different baselines with both CNN-based and ViT-based backbones, demonstrating its versatility and effectiveness across model architectures.
@@ 5.4 Robustness across Other Ground View Variations
Robustness across Locations (Table 7). In cross-view geo-localization, the VIGOR dataset is particularly challenging, because the two views are not center-aligned. In particular, VIGOR’s cross-area subset is regarded as a standard benchmark to test the model’s robustness to data across different locations, since training and test images are taken from different areas [44, 34, 5]. As shown in Table 7, ConGeo outperforms the baselines on both the cross-area and same-area subsets. Note that we report the performance of a single model for all FoVs, for ConGeo and Sample4Geo (grey background). The results indicate that ConGeo consistently improves the model’s robustness on cross-area data.
@@
Table 7: Comparison on University-1652 [41]. denotes the reproduced model since the pre-trained weights are not provided.
Methods Random FoVs Random Zooming Gaussian Noise Motion Blur Avg.
R@1 R@5 R@10 R@1% R@1 R@5 R@10 R@1% R@1 R@5 R@10 R@1% R@1 R@5 R@10 R@1% R@1
Sample4Geo 15.7 25.3 31.2 51.7 48.2 60.8 65.8 80.1 42.4 60.9 67.0 81.6 31.4 38.6 42.0 52.9 34.4
Sample4Geo + DA 85.5 95.1 96.6 98.8 44.5 60.6 66.7 82.7 0.2 1.0 1.7 8.8 16.8 23.2 25.9 36.0 36.8
ConGeo 84.2 95.1 97.0 99.4 68.7 80.3 83.8 92.1 45.8 64.4 70.3 83.4 32.5 40.7 43.5 53.5 57.8
Table 8: Comparison on unseen ground view variations between ConGeo and baselines on the CVUSA dataset. Random FoVs are between 0 to 360, Random Zooming is performed with a ratio between 0.5 and 2.0, Gaussian Noise [1], and Motion Blur [1] are added with severity: 5.
Robustness to Street Images (Table 7). Street images in University-1652 are typical examples of real-world limited FoV images. The St2S and S2St settings pose significant challenges as the limited FoV images can be captured from diverse locations with unknown orientations. Different from the experiments considering North-aligned panoramas, here we randomly sample two street view images from one location as the inputs to conduct single-view contrastive loss. Detailed descriptions can be found in the supplementary materials. Results in Table 7 show that the performance on both settings is improved.
Robustness to Unseen Ground View Variations (Table 8). In real-world scenarios, the variations of the ground view image are more diverse than orientation and FoV variations. In Table 8, we evaluate the model across four unseen ground view variations on the CVUSA dataset: Random FoVs, Random Zooming, Gaussian Noise, and Motion Blur. The ConGeo-augmented model exhibits comparable performance to the model with data augmentation on Random FoVs test, while demonstrating significant advantages over the two baselines when tested on other unseen ground view variations. According to Table 3 and 8, the improvements that task-oriented data augmentation brings are not transferable nor generalizable, on the contrary, ConGeo represents a way of unleashing the potential of data augmentation by enforcing the learning of invariances.
@@
Figure 5: ConGeo’s activation areas are more consistent across ground view variants. The Grad-CAM activation maps of the base model and ConGeo on the North-aligned and the unknown orientation setting. The orange box indicates the same area in different ground view variants.
Figure 6: ConGeo focuses on semantically consistent features for the task and its activated regions are consistent across view variation. Left: GradCam activations of the base model and ConGeo (green background) on the North-aligned and FoV = 180 settings. Right: focus on the regions highlighted by the coloured boxes.
@@ 7 Limitations
We showed that ConGeo leads to significant improvements when facing arbitrarily oriented ground images and diverse FoVs, except in the North-aligned setting. This performance drop is nearly unavoidable, as ConGeo tends to disrupt the over-reliance on spatial correspondence shortcuts. However, by keeping the original learning objectives of the base model (vanilla loss), ConGeo achieves competitive performance when orientation information is available (Table 2), and significantly improves the robustness when it is unknown. Additionally, we focus on ground view orientation and FoV variations, but other variations can be envisaged (e.g., zoom, color intensity, blur). We show ConGeo’s robustness to some of these variations unseen during training (Table 8). Finally, besides contrastive learning, other ways of aligning modalities (e.g., redundancy reduction) could also be considered. We discuss this in the supplementary materials.
@@ 8 Conclusion
Tackling situations where the ground view image orientation is unknown, or the FoV is limited, is crucial for real-world applications of cross-view image geo-localization. We propose ConGeo, a single- and cross-view contrastive method that enhances a model’s robustness to ground view variations by aligning image variations with their original representations. Experiments on four datasets and activation map analysis demonstrate that ConGeo consistently boosts the performance of state-of-the-art methods by a large margin when facing diverse challenging settings. With its adaptability to different base models and versatility to accommodate diverse ground view variations, ConGeo strives to be a step towards widening the applicability of geo-localization methods to the real world.
@@ Acknowledgments
We thank the anonymous reviewers for their constructive and thoughtful comments. We thank Haoyuan Li, Zimin Xia, Gencer Sümbül, Silin Gao, Valérie Zermatten, Zeming Chen, Tianqing Fang, Gaston Lenczner, Kuangyi Chen, Robin Zbinden, Giacomo May, Riccardo Ricci, Emanuele Dalsasso, and Sepideh Mamooler for providing helpful feedback on earlier versions of this work. We acknowledge the support from the CSC and EPFL Science Seed Fund and the support in part by the National Natural Science Foundation of China (NSFC) under Grant 62271355. AB gratefully acknowledges the support of the Swiss National Science Foundation (No. 215390), Innosuisse (PFFS-21-29), the EPFL Center for Imaging, Sony Group Corporation, and the Allen Institute for AI.
Supplementray Materials
@@
  • Additional experiments: architecture analysis (CNN-based and ViT-based encoders), transfer experiments, polar transformation, different paradigms of aligning modalities (Section S5).
@@
  • Visualization: additional samples and failure case analysis (Section S7).
@@
Appendix S3 Detailed Method Description
@@
Training Phase: In each mini-batch, we take paired ground-aerial images ($I_{Q}$, $I_{R}$) as input, and perform defined transformations $T_{q}$ (i.e., random shift, FoV cropping) to each ground view image $I_{q}$ ($I_{q}\in I_{Q}$), obtaining the transformed ground image $I_{q}{{}^{*}}=T_{q}(I_{q}|\theta,\alpha)$. We also perform augmentations in the base model to $I_{R}$ (described in the data preprocessing of Section S2), obtaining $I_{R}{{}^{*}}$. Then, the image batchs $I_{Q}$, $I_{Q}{{}^{*}}$, $I_{R}$ and $I_{R}{{}^{*}}$ are encoded by two encoders, a query encoder $E_{q}$ for encoding $I_{Q}$, $I_{Q}{{}^{*}}$, and a reference encoder $E_{r}$ for encoding $I_{R}$, $I_{R}{{}^{*}}$, with output feature embeddings $Q$, $Q{{}^{*}}$, $R$, and $R{{}^{*}}$ respectively. The optimization goal contains four targets, two single-view contrastive objectives, one for aligning the feature space between $Q$ and $Q{{}^{*}}$ while another for aligning the feature space between $R$ and $R{{}^{*}}$, a cross-view contrastive objective aligning the feature space between $R$ and $Q{{}^{*}}$, and a vanilla contrastive objective aligning the feature space between $R$ and $Q$. The total loss is a weighted summation of the four loss functions, where the ablation of weights will be studied in the next section.
@@
Street images in the University-1652 dataset are ground view images with unknown orientations and limited FoVs. In order to adapt ConGeo to street images without collecting a full-view panorama, we provide an alternative way of building the ground view single-view contrastive objective (Eq. (1) in the paper). Instead of enforcing the representations of ground view variants to be closer to the original ones, we construct the single-view contrastive objective by emphasizing the proximity between street images obtained from the same geographic locations but depicting different perspectives or angles. Since the street-to-satellite retrieval is a many-to-one matching, there is more than one street view image located in the region of satellite images. Therefore, for each street view sample, we randomly select another street image with the same location to construct the single-view contrastive objective for training.
Appendix S4 Hyper-parameter Analysis
@@
In order to achieve robustness across ground view variations, ConGeo applies a set of transformations on ground view images during training. The transformation includes shifting the ground view images with a random orientation angle $\theta$ and cropping the shifted images with a FoV angle $\alpha$. In the main experiments, we empirically set $\alpha$ to 180. Here, we provide detailed results by using different training FoV angles. Results are shown in Table S2. In the table, we report models trained with FoV images from 90, 180, 360, and random FoV from 0 to 360 on the unknown orientation and FoV settings. It can be seen that ConGeo’s performance is robust under different training FoVs, which consistently surpasses the baseline with targeted data augmentation by a large margin.
Method($\alpha$) FoV=360 FoV=180 FoV=90 FoV=70 Avg.
R@1 R@5 R@10 R@1% R@1 R@5 R@10 R@1% R@1 R@5 R@10 R@1% R@1 R@5 R@10 R@1% R@1
Sample4Geo+DA 84.1 95.0 96.9 99.4 63.6 84.6 90.1 98.2 32.2 55.1 64.5 87.4 21.5 42.0 51.6 79.9 50.4
ConGeo(90) 81.4 93.0 95.3 98.6 92.2 98.1 98.9 99.7 55.5 75.4 81.5 93.9 35.5 54.8 61.9 81.0 66.2
ConGeo(180) 85.2 95.1 96.9 98.9 92.3 97.9 98.7 99.7 55.9 73.2 79.0 90.9 37.1 55.7 62.8 81.4 67.6
ConGeo(360) 96.6 98.9 99.2 99.7 83.8 94.2 96.1 98.8 38.2 58.2 65.2 82.0 19.5 35.9 43.8 66.5 59.5
ConGeo(0-360) 63.5 80.0 85.4 95.5 82.8 94.1 96.3 99.3 55.2 77.3 83.0 94.7 46.4 67.1 75.2 91.6 62.0
Table S2: Analysis of training ConGeo with different FoV angle $\alpha$: 90, 180, 360, and random angles between 0-360, on the CVUSA dataset. Sample4Geo+DA means training Sample4Geo with targeted data augmentation (random shift and random FoVs 70-360).
Loss weights North-aligned FoV=90
@@
Table S3: Analysis of loss weights of single- and cross-view contrastive objectives on the CVUSA dataset.
LR North-aligned FoV=90